OpenStudy (anonymous):

Find the roots of z^3 = -27i thanks

OpenStudy (haseeb96):

z=-3 multiply with-1sqrt-1 z= -3i is the answer of it

OpenStudy (haseeb96):

and it is imaginary

OpenStudy (anonymous):

It's helpful to know that the powers of i form a cycle: i^1 = i I^2 = -1 i^3 = -i i^4 = 1 After this, it goes back around to i. Larger powers can be divided by 4 and you can just take the remainder to simplify powers of i. This should help with the -i part; the 27 should be straightforward enough.

OpenStudy (anonymous):

\[z^3=-27i\]You know that \(i^2=-1\) so we can write\[z^3=27i^3\]\[z^3-27i^3=0\]\[z^3-(3i)^3=0\]\[(z-3i)(z^2+3iz+9i^2)=0\]so u have\[z-3i=0\]and a quadratic equation\[z^2+3iz-9=0\]